OpenStudy (rayford):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) x + 6y = 1 2x + 5y = 9

OpenStudy (inkyvoyd):

Hey, do you have any idea how you would begin to solve this?

OpenStudy (rayford):

dude i have no clue whatsoever how to do this..

OpenStudy (inkyvoyd):

In that case. Let's solve a completely related problem!

OpenStudy (inkyvoyd):

5x + y = 1 -3x + 9y = 9 Okay... So we are supposed to use substitution right?

OpenStudy (rayford):

yes

OpenStudy (inkyvoyd):

btw, I'm not doing ur problem, cause I"m afraid I"m cheating on a test for you. THat would be bad, wouldn't it ;)

OpenStudy (rayford):

but ur not! my friend has this very complicated worksheet and he has been stuck on this problem for hours! so i wanted to surprise him with the answer..

OpenStudy (inkyvoyd):

5x + y = 1 -3x + 9y = 9 Now, we want to find an equation where we move a variable to one side, and "substitute" it in.

OpenStudy (inkyvoyd):

I don't want to help your "friend" cheat on his worksheet :)

OpenStudy (rayford):

ok..welll at least help me please?

OpenStudy (inkyvoyd):

Sure. That's what I'm doing this example question for. I can do as many of these as I want and I won't be cheating or helping people cheat :)

OpenStudy (rayford):

ok well continue please?

OpenStudy (inkyvoyd):

5x + y = 1 ->call this eq1 -3x + 9y = 9 ->call this eq2

OpenStudy (inkyvoyd):

eq1: 5x+y=1 -> y=1-5x

OpenStudy (inkyvoyd):

eq2: -3x+9y=9 -3x+9(y)=9

OpenStudy (inkyvoyd):

But, we know from eq1 that (y)=1-5x

OpenStudy (inkyvoyd):

so, -3x+9(1-5x)=9 This is the biggest part. Tell me if it makes sense or not.

OpenStudy (rayford):

it does

OpenStudy (inkyvoyd):

I got rid of the y. Now there are just x's. Why don't we solve that now.

OpenStudy (inkyvoyd):

-3x+9(1-5x)=9 -3x+9-45x=9 -48x=0 x=0

OpenStudy (inkyvoyd):

BUT WHAT DOES Y EQUAL?

OpenStudy (inkyvoyd):

5(x) + y = 1 -> eq1 -3(x) + 9y = 9 -> eq2 we know x equals 0...

OpenStudy (inkyvoyd):

Working with either equation at this point to find y works. I'll just do eq1 5(x)+y=1 but x=0 5(0)+y=1 0+y=1 y=1

OpenStudy (inkyvoyd):

In other words, x=0, y=1

OpenStudy (rayford):

hmmmm ok, ok let me look over this..

OpenStudy (inkyvoyd):

Look over it long and hard, and when you are ready, let's do the actual test problem :)

OpenStudy (inkyvoyd):

Actually, you'll be doing most of the work, while I nudge you along. I don't like getting banned by moderators for helping people cheat :P

OpenStudy (rayford):

im so sorry i fell asleep. but ended up figuring it out thank you!